How much money is invested in bonds and how much money is invested in certificates if the municipal bonds yeilding 6% annually and certificates of deposit yeilding 8% annually? The amount of investments is $24,636 and the annual income is 1769.20.
Answer by stanbon(75887)

How much money is invested in bonds and how much money is invested in certificates if the municipal bonds yeilding 6% annually and certificates of deposit yeilding 8% annually? The amount of investments is $24,636 and the annual income is 1769.20.
---------------
Quantity Deposited:: s + e = 24,636
Interest Earned::: 6s +8e = 176920
-------------------------
Modify for elimination::
3s + 3e = 3*24636
3s + 4e = 88460
---------------------
Subtract and solve for "e":
e = $14,552 (amt. invested at 8%)
---
s = 24636-14552 = $10,084 (amt. invested at 6%)
=================
Cheers,
Stan H.
=================